Bookkeeping of the Calorimeter Commissioning Data
This page provides useful information about the data samples stored during the commissioning phase of the calorimeters.
Much more data files than described here have be stored. Only the most relevant are detailled below.
If you are looking for other data please
send me a mail
ADDITIONAL INFORMATION ABOUT DATA TAKING DURING COMMISSIONING SESSIONS
CAN BE FOUND
IN THE LOG BOOK
Instructions
The DAQ commissioning data possibly exist in several format : mdf and/or raw depending on the DAQ setup :
Each event in mdf format contains a single trigger. Each event in raw format possibily matches a sequence of several consecutive triggers. In this case the corresponding rawEvents are located in different TES pathes : DAQ/Prev1, DAQ/Next2, ...
The 'Time Alignment Events' (TAE) correspond to a trigger sequence of consecutive BCID.
OBSOLETE : MDF FORMAT NOW SUPPORTS TAE STRUCTURE. 'RAW' FILES ARE NO LONGER PRODUCED
data location
Data taken using the LHCb storage system
The stored data have a
MDF format (with .raw extension).
On the online system the data files are stored at location like :
/daqarea/lhcb/data/2008/RAW/LHCb/PHYSICS/[run]/
The data are also transfered on Castor at :
/castor/cern.ch/ grid /lhcb/data/2008/RAW/LHCb/PHYSICS/
(some old files had been stored at /castor/cern.ch/lhcb/online/point8/2008/RAW/LHCb/PHYSICS/)
Data taken using the local Event Builder (within Orwell framework) OBSOLETE
The data files can be found Castor storage system at :
/castor/cern.ch/lhcb/online/point8/calo/
Some of the files are also replicated at :
/afs/cern.ch/lhcb/group/calo/ecal/vol9/odescham/commissioning/data/
Most of the data listed below have been analysed using the
Orwell application and monitoring ntuples have been produced. These ntuples (root and hbook format) can be found at :
/castor/cern.ch/lhcb/online/point8/calo/ntp/
and are partially replicated at :
/afs/cern.ch/lhcb/group/calo/ecal/vol9/odescham/commissioning/ntp/
Naming convention
Data taken using the LHCb storage system
The file name convention is : [run]_000XXXX.raw
Data taken with the local Event Builder (Orwell framework) OBSOLETE
The convention for data file naming is :
[data]_[date]_[number].[type]
with
[data] = hcal / prs / ecal / ...
[type] = mdf / raw / root / hbook
[date] format is [mm][dd]
The files are usually located in directories named
com[date]/ with [mm][yy] as the date format.
---
August 2008
Data [LHCb continuous shifts (cosmics run)]
run |
#sequences |
Triggger type |
rate |
TAE window |
comment |
32841 |
314k |
TCK=0x0002 |
10Hz |
+-5 |
log , log |
32586 |
?? |
TCK=0x0009 |
<0.1Hz |
+-5 |
log , log CONTAIN TED EVENTS |
32583 |
?? |
TCK=0x0009 |
<0.1Hz |
+-5 |
log , log CONTAIN TED EVENTS |
32578 |
?? |
TCK=0x0009 |
<0.1Hz |
+-5 |
log CONTAIN TED EVENTS |
32576 |
?? |
TCK=0x0009 |
<0.1Hz |
+-5 |
log CONTAIN TED EVENTS |
32501 |
?? |
TCK=0x0009 |
<0.1Hz |
+-5 |
log CONTAIN TED EVENTS |
32500 |
?? |
TCK=0x0009 |
<0.1Hz |
+-5 |
log CONTAIN TED EVENTS |
32498 |
?? |
TCK=0x0009 |
<0.1Hz |
+-5 |
log CONTAIN TED EVENTS |
32493 |
114 |
TCK=0x0009 |
<0.1Hz |
+-5 |
log CONTAIN TED EVENTS |
32484 |
215 |
TCK=0x0009 |
<0.1Hz |
+-5 |
log CONTAIN TED EVENTS |
32481 |
51 |
TCK=0x0009 |
<0.1Hz |
+-5 |
log CONTAIN TED EVENTS |
32479 |
?? |
TCK=0x0009 |
<0.1Hz |
+-5 |
log CONTAIN TED EVENTS |
32477 |
?? |
TCK=0x0009 |
<0.1Hz |
+-5 |
log CONTAIN TED EVENTS |
32474 |
?? |
TCK=0x0009 |
<0.1Hz |
+-5 |
log CONTAIN TED EVENTS |
32471 |
?? |
TCK=0x0009 |
<0.1Hz |
+-5 |
log |
32468 |
?? |
TCK=0x0009 |
<0.1Hz |
+-5 |
log |
32457 |
?? |
TCK=0x0009 |
<0.1Hz |
+-5 |
log |
32445 |
?? |
TCK=0x0009 |
1Hz |
+-5 |
log |
32443 |
?? |
TCK=0x0002 |
10Hz |
+-5 |
log |
32442 |
11k |
TCK=0x0002 |
10Hz |
+-5 |
log |
32439 |
5k |
TCK=0x0002 |
10Hz |
+-5 |
log |
32434 |
? |
TCK=0x0002 |
10Hz |
+-5 |
log |
32432 |
? |
TCK=0x0002 |
10Hz |
+-5 |
log |
32428 |
32k |
TCK=0x0002 |
10Hz |
+-5 |
log |
32406 |
50k |
TCK=0x0002 |
10Hz |
+-5 |
log , log |
32385 |
145k |
TCK=0x0002 |
10Hz |
+-5 |
log , log |
32381 |
?? |
TCK=0x0002 |
10Hz |
+-5 |
log , log |
32380 |
?? |
TCK=0x0002 |
10Hz |
+-5 |
log , log |
32033 |
?? |
TCK=0x0002 |
10Hz |
+-5 |
log |
32026 |
??? |
TCK=0x0002 |
10Hz |
+-5 |
log , log |
32025 |
7k |
TCK=0x0002 |
10Hz |
+-5 |
log , log |
32024 |
??? |
TCK=0x0002 |
10Hz |
+-5 |
log , log |
32023 |
5k |
TCK=0x0002 |
10Hz |
+-5 |
log , log |
32022 |
8k |
TCK=0x0002 |
10Hz |
+-5 |
log , log |
32021 |
6k |
TCK=0x0002 |
10Hz |
+-5 |
log , log |
32020 |
20k |
TCK=0x0002 |
10Hz |
+-5 |
log , log |
32011 |
90k |
TCK=0x0002 |
10Hz |
+-5 |
log , log |
32009 |
15k |
TCK=0x0002 |
10Hz |
+-5 |
log , log |
32007 |
20k |
TCK=0x0002 |
10Hz |
+-5 |
log , log |
32005 |
48k |
TCK=0x0002 |
10Hz |
+-5 |
log , log |
32001 |
26k |
TCK=0x0002 |
10Hz |
+-5 |
log |
31874 |
126k |
TCK=0x0002 |
10Hz (+1Hz calib) |
+-5 |
log , log |
31856 |
??? |
TCK=0x0002 |
10Hz (+1Hz calib) |
+-5 |
log , log |
31855 |
??? |
TCK=0x0002 |
10Hz (+1Hz calib) |
+-5 |
log , log |
31848 |
??? |
TCK=0x0002 |
10Hz (+1Hz calib) |
+-5 |
log , log |
31846 |
74k |
TCK=0x0002 |
10Hz (+1Hz calib) |
+-5 |
log , log |
31845 |
2k |
TCK=0x0002 |
10Hz (+1Hz calib) |
+-5 |
log , log |
31843 |
15k |
TCK=0x0002 |
10Hz (+1Hz calib) |
+-5 |
log , log |
31842 |
1.3k |
TCK=0x0002 |
10Hz (+1Hz calib) |
+-5 |
log , log |
31841 |
89k |
TCK=0x0002 |
10Hz (+1Hz calib) |
+-5 |
log , log |
31838 |
33k |
TCK=0x0002 |
10Hz (+1Hz calib) |
+-5 |
log ,log |
31821 |
78k |
TCK=0x0002 |
10Hz (+1Hz calib) |
+-5 |
log |
31820 |
8k |
TCK=0x0002 |
10Hz (+1Hz calib) |
+-5 |
log |
31817 |
11k |
TCK=0x0002 |
10Hz (+1Hz calib) |
+-5 |
log |
31814 |
?? |
TCK=0x0002 |
10Hz (+1Hz calib) |
+-5 |
log |
31813 |
?? |
TCK=0x0002 |
10Hz (+1Hz calib) |
+-5 |
log |
31811 |
?? |
TCK=0x0002 |
10Hz (+1Hz calib) |
+-5 |
log |
31810 |
?? |
TCK=0x0002 |
10Hz (+1Hz calib) |
+-5 |
log |
31809 |
?? |
TCK=0x0002 |
10Hz (+1Hz calib) |
+-5 |
log |
31808 |
?? |
TCK=0x0002 |
10Hz (+1Hz calib) |
+-5 |
log |
31806 |
?? |
TCK=0x0002 |
10Hz (+1Hz calib) |
+-5 |
log |
31797 |
?? |
TCK=0x0002 |
10Hz (+1Hz calib) |
+-3 |
log |
31794 |
?? |
TCK=0x0002 |
10Hz (+1Hz calib) |
+-3 |
log , log |
31793 |
13k |
TCK=0x0002 |
10Hz |
+-3 |
log |
31792 |
90k |
TCK=0x0002 |
10Hz |
+-3 |
log |
31789 |
?? |
TCK=0x0002 |
10Hz |
+-3 |
log |
31739 |
?? |
TCK=0x0002 |
10Hz |
+-2 |
log |
31734 |
37k |
TCK=0x0002 |
10Hz |
+-2 |
log |
31732 |
?? |
TCK=0x0002 |
10Hz |
+-2 |
log ,log |
31731 |
22k |
TCK=0x0002 |
10Hz |
+-2 |
log |
31727 |
9k |
TCK=0x0002 |
10Hz |
+-2 |
log |
31725 |
12k |
TCK=0x0002 |
10Hz |
+-2 |
log |
31723 |
75k |
TCK=0x0002 |
10Hz |
+-2 |
log |
31720 |
36k |
TCK=0x0002 |
10Hz |
+-2 |
log |
31716 |
2k |
TCK=0x0002 |
10Hz |
+-2 |
log |
31715 |
8k |
TCK=0x0002 |
10Hz |
+-2 |
log |
31714 |
23k |
TCK=0x0002 |
10Hz |
+-2 |
log |
31557 |
180k |
TCK=0x0002 |
10Hz |
+-2 |
log |
31551 |
?? |
TCK=0x000D |
??Hz |
+-2 |
log |
31550 |
?? |
TCK=0x000C |
??Hz |
+-2 |
log |
31549 |
?? |
TCK=0x000C |
??Hz |
+-2 |
log |
31558 |
?? |
TCK=0x000C |
??Hz |
+-2 |
log |
31227 |
150k |
TCK=0x0002 |
10Hz |
+-2 |
log |
31225 |
150k |
TCK=0x0002 |
10Hz |
+-2 |
log , log |
31216 |
8k |
TCK=0x0002 |
10Hz |
+-2 |
log |
31160 |
25k |
TCK=0x0002 |
10Hz |
+-2 |
log |
31060 |
?? |
TCK=0x0002 |
10Hz |
?? |
log |
30942 |
?? |
TCK=0x0009 |
??Hz |
+-7 |
log |
30947 |
?? |
TCK=0x0009 |
??Hz |
+-7 |
log |
30915 |
?? |
TCK=0x0009 |
0.1Hz |
+-7 |
log |
30800 |
?? |
TCK=0x0003 |
??Hz |
+-5 |
log |
30799 |
?? |
TCK=0x0009 |
0.1Hz |
+-5 |
log |
30787 |
?? |
TCK=0x0009 |
0.1Hz |
+-7 |
log |
30767 |
?? |
TCK=0x0009 |
0.1Hz |
+-7 |
log CONTAIN TED TI8 INJECTION EVENTS |
30764 |
?? |
TCK=0x0009 |
0.1Hz |
+-7 |
log CONTAIN TED TI8 INJECTION EVENTS |
30763 |
?? |
TCK=0x0009 |
0.1Hz |
+-7 |
|
30762 |
?? |
TCK=0x0009 |
0.1Hz |
+-7 |
log |
30756 |
?? |
TCK=0x0009 |
0.1Hz |
+-7 |
log |
30664 |
34.5k |
TCK=0x0002 |
10Hz |
+-2 |
log |
30662 |
24k |
TCK=0x0002 |
10Hz |
+-2 |
log |
30661 |
18k |
TCK=0x0002 |
10Hz |
+-2 |
log |
30660 |
?? |
TCK=0x0002 |
10Hz |
+-2 |
log |
30659 |
7.5k |
TCK=0x0002 |
10Hz |
+-2 |
log |
30654 |
?? |
TCK=0x0002 |
10Hz |
+-2 |
log |
30653 |
7.5k |
TCK=0x0002 |
10Hz |
+-2 |
log |
30658 |
23k |
TCK=0x0002 |
10Hz |
+-2 |
log |
30645 |
6k |
TCK=0x0002 |
10Hz |
+-2 |
log |
30641 |
15k |
TCK=0x0002 |
10Hz |
+-2 |
log |
30640 |
2k |
TCK=0x0002 |
10Hz |
+-5 |
log |
30639 |
13k |
TCK=0x0002 |
10Hz |
+-2 |
log |
30440 |
?? |
TCK=0x0002 |
10Hz |
+-4 |
log |
30436 |
?? |
TCK=0x0002 |
10Hz |
+-4 |
log ,log |
30434 |
41k |
TCK=0x0002 |
10Hz |
+-4 |
log ,log |
30433 |
5.5k |
TCK=0x0002 |
10Hz |
+-2 |
log |
30432 |
2.5k |
TCK=0x0002 |
10Hz |
+-2 |
log |
30431 |
?? |
TCK=0x0002 |
10Hz |
+-2 |
log |
30420 |
62.5k |
TCK=0x0002 |
10Hz |
+-2 |
log |
30415 |
13k |
TCK=0x0002 |
10Hz |
+-2 |
log |
30406 |
17k |
TCK=0x0002 |
10Hz |
+-2 |
log |
Data (17/08/2008) [Ecal LED]
- data : Ecal
- settings :
- scan setup : Each LED is flashed with 1/8 trigger frequency - LEDTSB settings: default
- run coord. : Irina
run |
mdf/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
29970 |
x0 |
20k |
LED |
calibration |
56Hz |
5(TAE) |
nominal HV; nominal LED intensity |
29972 |
x0 |
20k |
LED |
calibration |
56Hz |
5(TAE) |
PM Gain 50K; LED intensity 0.3V |
29973 |
x0 |
20k |
LED |
calibration |
56Hz |
5(TAE) |
PM Gain 50K; LED intensity 0.4V |
29974 |
x0 |
20k |
LED |
calibration |
56Hz |
5(TAE) |
PM Gain 25K; LED intensity 0.4V |
29975 |
x0 |
20k |
LED |
calibration |
56Hz |
5(TAE) |
PM Gain 15K; LED intensity 0.4V |
29976 |
x0 |
20k |
LED |
calibration |
56Hz |
5(TAE) |
PM Gain 15K; LED intensity 0.6V |
29977 |
x0 |
20k |
LED |
calibration |
56Hz |
5(TAE) |
PM Gain 5K; LED intensity 0.6V |
29978 |
x0 |
100k |
LED |
calibration |
5Hz |
5(TAE) |
PM gain nominal; LED intensity nominal |
Data (8-10/08/2008) [Ecal LED]
- data : Ecal
- settings :
- scan setup : Each LED is flashed with 1/8 trigger frequency - LEDTSB settings: default
- run coord. : Irina
run |
mdf/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
29358 |
x0 |
20k |
LED |
calibration |
56Hz |
5(TAE) |
PM gain 100K; LED intensity = 0 |
29323 |
x0 |
100k |
LED |
calibration |
4Hz |
5(TAE) |
PM gain nominal; LED intensity nominal |
July/August 2008
Data [LHCb commissioning shifts (cosmics run)]
July 2008
Data (18 & 20 /07/2008) [Ecal LED]
- data : Ecal
- settings :
LEDTSB_del_A= ecA_TSBrecipe_24893
LEDTSB_del_C = ecC_TSBrecipe_26653
- scan setup : Each LED is flashed with 1/8 trigger frequency
- run coord. : Irina / Iouri / Marie-Hélène
run |
mdf/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
27308 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
"cosmic" PM setting, LED_intensity=0 |
|
27324 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
alternative PM setting (1), LED_LV=nominal |
27327 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
alternative PM setting (1), LED_LV=nominal-0.05V |
27328 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
alternative PM setting (1), LED_LV=nominal-0.1V |
27329 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
alternative PM setting (1), LED_LV=nominal+0.05V |
27330 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
alternative PM setting (1), LED_LV=nominal+0.1V |
|
27341 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal |
27346 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal-0.05V |
27349 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal-0.10V |
27354 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal-0.15V |
27361 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal-0.20V |
27364 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal-0.25V |
27368 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal-0.30V |
|
27370 |
x0 |
10.4k |
LED scan (2) |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal-0.30V |
|
27376 |
x0 |
100k |
LED |
calibration |
2Hz |
5(TAE) |
stability run - nominal PM setting, LED_LV=nominal-0.30V |
|
27402 |
x0 |
25k |
LED (3) |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal |
27403 |
x0 |
10.4k |
LED scan (2) |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal |
27404 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal |
|
27405 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal-0.05V |
27406 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal-0.10V |
27407 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal-0.15V |
27408 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal-0.20V |
27409 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal-0.30V |
27410 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal-0.40V |
27411 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal+0.05V |
27412 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal+0.10V |
27413 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal+0.15V |
|
27414 |
x0 |
25k |
LED (4) |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal |
|
27427 |
x0 |
25k |
LED (5) |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal |
27428 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal-0.05V |
27429 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal-0.10V |
27430 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal-0.15V |
27431 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal-0.20V |
27432 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal+0.05V |
27433 |
x0 |
25k |
LED |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal+0.10V |
27434 |
x0 |
20.8k |
LED scan (6) |
calibration |
56Hz |
5(TAE) |
nominal PM setting, LED_LV=nominal |
- (1) PM gain is "nominal", but corrected on the spread of PM quantum efficiencies + different
formula to recalculate HV on PM ("alternative" setting which is not stored in DB); LED LV = nominal
- (2) Time scan, 13 points, 800 evt/point
- (3) 20/07/2008 : LED LV settings are changed in DB regarding previous data taken on 18/07/2008.
- (4) Reference sample after switching OFF and ON HV
- (5) WRONG LED setting discovered for run 27402 to 27414. They have been repeated with correct setting with run 27427 to 27434
- (6) Time scan, 13 points, 1600 evt/point
Data (11/07/2008) [Ecal LED scan]
- data : Ecal A+C
- scan setup :
- 800 ev/step
- each PMT flashing by 1/8 trigger freq
- steps=13
- tot_ev= 10400
- Calibration trigger rate = 56 Hz
- settings :
- LEDTSB_del_C=150 ns
- LEDTSB_del_A=ecA_TSBrecipe_24893 (for run 25541 -> 26674)
- LEDTSB_del_A=ecA_TSBrecipe_23700 (for run 26775->26777)
- LED_intens=nominal cnt
- FErecipe-Default
- data location : /castor/cern.ch/grid/lhcb/data/2008/RAW/ HCAL/PHYSICS_TimeScan/[RUN]
- run coord. : Anatoli / Marie-Noelle
run |
mdf/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
26651 |
x0 |
10.4k |
LED scan |
calibration |
56Hz |
5(TAE) |
HV_A = nominal? HV_C = 0.5nominal, LED_intens=nominal |
26653 |
x0 |
10.4k |
LED scan |
calibration |
56Hz |
5(TAE) |
HV_A = nominal? HV_C = nominal, LED_intens=nominal |
26656 |
x0 |
24k |
LED NO scan |
calibration |
56Hz |
5(TAE) |
HV_A = nominal? HV_C = nominal, LED_intens=nominal |
26657 |
x0 |
24k |
LED NO scan |
calibration |
56Hz |
5(TAE) |
HV_A = nominal? HV_C = nominal, LED_intens=nominal-0.1V cnt |
26661 |
x0 |
24k |
LED NO scan |
calibration |
56Hz |
5(TAE) |
HV_A = nominal? HV_C = nominal, LED_intens=nominal-0.2V cnt |
26663 |
x0 |
24k |
LED NO scan |
calibration |
56Hz |
5(TAE) |
HV_A = nominal? HV_C = nominal, LED_intens=nominal-0.3V cnt |
26664 |
x0 |
10.4k |
LED scan |
calibration |
56Hz |
5(TAE) |
HV_A = nominal? HV_C = nominal, LED_intens=nominal-0.3V cnt |
26665 |
x0 |
24k |
LED scan |
calibration |
56Hz |
5(TAE) |
HV_A = nominal? HV_C = nominal, LED_intens=nominal-0.4V cnt |
26666 |
x0 |
24k |
LED scan |
calibration |
56Hz |
5(TAE) |
HV_A = nominal? HV_C = nominal, LED_intens=nominal-0.5V cnt |
26667 |
x0 |
24k |
LED scan |
calibration |
56Hz |
5(TAE) |
HV_A = nominal? HV_C = nominal, LED_intens=nominal-0.6V cnt |
26668 |
x0 |
24k |
LED scan |
calibration |
56Hz |
5(TAE) |
HV_A = nominal? HV_C = nominal, LED_intens=nominal-0.7V cnt |
26669 |
x0 |
10.4k |
LED scan |
calibration |
56Hz |
5(TAE) |
HV_A = nominal? HV_C = nominal, LED_intens=nominal-0.7V cnt |
26670 |
x0 |
24k |
LED scan |
calibration |
56Hz |
5(TAE) |
HV_A = nominal? HV_C = nominal, LED_intens=nominal-0.8V cnt |
26671 |
x0 |
24k |
LED scan |
calibration |
56Hz |
5(TAE) |
HV_A = nominal? HV_C = nominal, LED_intens=nominal-0.9V cnt |
26673 |
x0 |
24k |
LED scan |
calibration |
56Hz |
5(TAE) |
HV_A = nominal? HV_C = nominal, LED_intens=nominal-1.0V cnt |
26674 |
x0 |
2k |
LED NO scan |
calibration |
56Hz |
5(TAE) |
HV_A = nominal? HV_C = nominal, LED_intens=nominal-0.3V cnt |
26675 |
x0 |
2k |
LED NO scan |
calibration |
56Hz |
5(TAE) |
HV_A = nominal? HV_C = nominal, LED_intens=nominal-0.3V cnt |
26676 |
x0 |
2k |
LED NO scan |
calibration |
56Hz |
5(TAE) |
HV_A = nominal? HV_C = nominal, LED_intens=nominal-0.4V cnt |
26676 |
x0 |
10.4k |
LED scan |
calibration |
56Hz |
5(TAE) |
HV_A = nominal? HV_C = nominal, LED_intens=nominal-0.4V cnt |
26677 |
x0 |
100k |
LED NO scan |
calibration |
2Hz |
5(TAE) |
HV_A = nominal? HV_C = nominal, LED_intens=nominal-0.4V cnt |
Data (07/07/2008) [Ecal & Hcal LED scan]
- data : Ecal A+C OR Hcal A+C
- scan setup :
- 1600 ev/step
- each PMT flashing by 1/8 trigger freq
- steps=13
- tot_ev= 20800
- Calibration trigger rate = 56 Hz
- settings :
- HV = nominal
- LEDTSB_del_C=150 ns (for Hcal)
- LEDTSB_del_A=recipe_ecA_TSBrecipe_24893 (for Ecal) LEDTSB_del_A=150ns (for Hcal)
- LED_intens=nominal cnt
- FErecipe-Default
- data location : /castor/cern.ch/grid/lhcb/data/2008/RAW/ HCAL/PHYSICS_TimeScan/[RUN]
- run coord. : Anatoli / Marie-Noelle / Iouri
run |
mdf/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
26292 |
x0 |
20.8k |
LED scan |
calibration |
56Hz |
5(TAE) |
Ecal LEDTSB_del_C=150 ns |
26299 |
x0 |
20.8k |
LED scan |
calibration |
56Hz |
5(TAE) |
Ecal LEDTSB_del_C=ecC_TSBrecipe_23700 |
26308 |
x0 |
100k |
LED NO scan |
calibration |
56Hz |
5(TAE) |
Ecal LEDTSB_del_C=ecC_TSBrecipe_23700 |
26340 |
x0 |
96k |
LED NO scan |
calibration |
56Hz |
5(TAE) |
Ecal LEDTSB_del_C=ecC_TSBrecipe_23700 (1) |
26361 |
x0 |
20.8k |
LED scan |
calibration |
56Hz |
5(TAE) |
Hcal LED_intens=nominal cnt |
26362 |
x0 |
20.8k |
LED scan |
calibration |
56Hz |
5(TAE) |
Hcal LED_intens=0.95 nominal cnt |
26363 |
x0 |
100k |
LED NO scan |
calibration |
56Hz |
5(TAE) |
Hcal HV=Physics (Etmax~15GeV) |
(1) 10 min pause + 1600 event collect, time Points=60.
Started at 18:10 of 4.07.08, stopped at 4:33 of 5.07.08.
The run includes two raw files.
The LV monitoring started at18:00 and stopped 6:13.
Data (04/07/2008) [Hcal LED scan]
data : Hcal A+C
scan : 1600 ev/step, each PMT flashing by 1/8 trigger freq, steps=13, tot_ev= 20800, Calibration trigger rate = 56 Hz
setting : HV = nominal, LEDTSB_del_C=150 ns; LEDTSB_del_A= 150 ns ; FErecipe-Default.
data location : /castor/cern.ch/grid/lhcb/data/2008/RAW/ HCAL/PHYSICS_TimeScan/[RUN]
run |
mdf/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
26361 |
x0 |
21k |
LED |
calibration |
56Hz |
5(TAE) |
LED_intens=nominal cnt |
26362 |
x0 |
21k |
LED |
calibration |
56Hz |
5(TAE) |
LED_intens= 0.95 * nominal cnt (1) |
(1) the ODIN time is available in run 26362, where "internal UTC" was set for ODIN
Data (01-05/07/2008) [Global Commissioning cosmics runs]
Data content : all files contain at least : the whole calo (A+C for Spd/Prs/Ecal/Hcal) + L0Calo + L0DU + ODIN
Data are available in /daqarea/.../LHCb/PHYSICS/ except 26040 and 26046 which are in /LHCb/PHYSICS_COSMICS/
01/07/2008
run |
mdf/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
025941 (log) |
x0 |
?? |
cosmics |
L0DU (CaloCosmics) |
?? |
11(TAE) |
OTC + IT + Calo (HV on near the end of file) |
025950 |
x0 |
22k |
cosmics |
L0DU (CaloCosmics) |
?? |
11(TAE) |
data: Calo + OTC |
025951 |
x0 |
?? |
cosmics |
L0DU (CaloCosmics) |
?? |
11(TAE) |
data: Calo + OTC |
025953 |
x0 |
?? |
cosmics |
L0DU (CaloCosmics) |
?? |
11(TAE) |
data: Calo + OTC |
025956 |
x0 |
?? |
cosmics |
L0DU (CaloCosmics) |
?? |
11(TAE) |
data: Calo + OTC |
025975 |
x0 |
187k |
cosmics |
L0DU (CaloCosmics) |
?? |
11(TAE) |
data: Calo (night run : ~6h long) |
02/07/2008
'Standard' cosmics triggers
run |
mdf/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
026027 |
x0 |
?? |
cosmics |
L0DU (Calo OR Muon : TCK=0x0001) (1) |
60Hz |
5(TAE) |
data: Calo + MuonC +OTC + TT |
026040 (log) |
x0 |
21k |
cosmics |
L0DU (Calo) |
10Hz |
5(TAE) |
data: Calo + MuonC +OTC + TT + IT |
026046 (log) |
x0 |
147 |
cosmics |
L0DU (Calo OR Muon : TCK=0X0001) |
60Hz |
5(TAE) |
data: Calo + MuonC +OTC + TT +IT |
026066 |
x0 |
?? |
cosmics |
L0DU (Calo OR Muon : TCK=0x0001) |
60Hz |
5(TAE) |
data: Calo + MuonC +OTC + TT +IT |
026069 (log) |
x0 |
19k |
cosmics |
L0DU (Calo OR Muon : TCK=0x0001) |
60Hz |
7(TAE) |
data: Calo + MuonC +OTC + TT +IT |
026070 (log) |
x0 |
4k |
cosmics |
L0DU (Calo) |
60Hz |
5(TAE) |
data: Calo + MuonC +OTC + TT +IT |
026071 (log) |
x0 |
20k |
cosmics |
L0DU (Calo) |
60Hz |
5(TAE) |
data: Calo + MuonC +OTC + TT +IT |
(1) : step towards a complete
TCK (Trigger Configuration Key) mechanism for L0
From now the 'standard' cosmics algorithm
S1[Ecal>20 AND Hcal>5] OR S2[Muon1>0]
has L0-TCK=0x0001 in L0DU data bank.
Warning : all data taken before had
TCK=0 whatever the actual algorithm was !
'Special' cosmics triggers
run |
mdf/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
026078 |
x0 |
?? |
cosmics |
L0DU TCK=0x0005 (1) |
1Hz |
5(TAE) |
data: Calo + MuonC |
026079 (log) |
x0 |
?? |
cosmics |
L0DU TCK=0x0005 (1) |
1Hz |
5(TAE) |
data: Calo + MuonC |
026080 (log) |
x0 |
?? |
cosmics |
L0DU TCK=0x0006 (2) |
1Hz |
5(TAE) |
data: Calo + MuonC Warning : SpdMult has been time-aligned during the run (+1BX) |
026081 (log) |
x0 |
1k |
cosmics |
L0DU TCK=0x0007 (3) |
0.25Hz |
5(TAE) |
data: Calo + MuonC |
(1) TCK=0x0005 : S1=[Electron(Et)>20 AND Muon1(Pt)>0] OU S2=[Hadron(Et)>5 AND [Muon1(Pt)>0]
(2) TCK=0x0006 : S1=[Spd(Mult)>1]
(3) TCK=0x0007 : S1=[Spd(Mult)>0 AND Muon1(Pt)>0]
03/07/2008
run |
mdf/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
026194 (log) |
x0 |
?? |
cosmics |
L0DU (Calo OR Muon : TCK=0x0001) |
?? |
?? (TAE) |
data: Calo + MuonC +IT |
04/07/2008
run |
mdf/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
026227 (log) |
x0 |
58k |
cosmics |
L0DU (Calo OR Muon : TCK=0x0001) |
?? |
?? (TAE) |
data: Velo, TT, IT, OT (HV off), Calo and MuonC |
026231 (log) |
x0 |
44k |
cosmics |
L0DU (Calo OR Muon : TCK=0x0001) |
?? |
7(TAE) |
data: Velo, TT, IT, OTC (with HV), Calo, MuonC |
026246 (log) |
x0 |
14k |
cosmics |
L0DU (Calo OR Muon : TCK=0x0001) |
?? |
5(TAE) |
data: All detectors but RICH |
June 2008
Data (20/06/2008) [Ecal-A LED]
Data content : Ecal-A
ECAL data (1600 ev/step, each PMT flashing by 1/8 trigger freq, steps=13, tot_ev= 20800,
Calibration trigger rate = 56 Hz).
Data are available at : /castor/cern.ch/grid/lhcb/data/2008/RAW/ECAL/PHYSICS_TimeScan/
HV = nominal, LEDTSB_del=calculated recipe; LED_intens=nominal cnt; FErecipe-Default
run |
mdf/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
024893 |
x0 |
20800 |
LED scan |
calibration |
56Hz |
5(TAE) |
|
024902 |
x0 |
20800 |
LED scan |
calibration |
56Hz |
5(TAE) |
Collection time : 62 mn |
025039 |
x0 |
20800 |
LED scan |
calibration |
56Hz |
5(TAE) |
FE_PIN corrected latency L=151 (FE14 ch_1-16), L=150 (FE15 ch_1-16) |
025041 |
x0 |
125k |
LED no-scan (1) |
calibration |
56Hz |
5(TAE) |
FE_PIN corrected latency L=151 (FE14 ch_1-16), L=150 (FE15 ch_1-16) |
(1) : without scan for stability
1600ev/point, 80 points.with 10 min wait between points ; 125k events
started 22.06.08 at 16:15 and finished at (3:38) 6:27 of 23.06.2008
Data (06/06/2008) [Hcal-A LED]
Data content :
- Hcal-A
- long run (13h43mn) for stability study
- run structure (1600 evts @ 56 Hz+ 5mn waiting)*150 step
- each PMT flashing with 1/8 of the trigger rate
- HV nominal - LEDTSB_del=150ns - LED intensity : nominal
- Default FERecipe with corrected time setting for PIN
- data available on /HCAL/PHYSICS_TimeScan/24432
run |
mdf/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
024432 |
x0 |
240k |
LED |
calibtation |
56Hz |
5(TAE) |
|
Data (02-06/06/2008) [Global Commissioning cosmic runs]
Data content :
- CALO A+C-sides + ODIN
- run coord: Olivier C.
run |
mdf/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
023776 (mond.) |
x0 |
34k |
cosmics |
L0DU (CaloCosmics or Muon) |
?? |
5(TAE) |
|
023837 (tues.) |
x0 |
6k |
cosmics |
L0DU (CaloCosmics or Muon) |
(5+1)Hz |
5(TAE) |
(log_L0Trigger |
023974 (wedn.) |
x0 |
?? |
cosmics |
L0DU (CaloCosmics) |
~10 Hz |
5(TAE) |
(1) |
023985 (wedn.) |
x0 |
?? |
cosmics |
L0DU (CaloCosmics) |
~10 Hz |
5(TAE) |
(1) |
023989 (wedn.) |
x0 |
4k |
cosmics |
L0DU (CaloCosmics) |
~10 Hz |
5(TAE) |
(1) |
023990 (wedn.) |
x0 |
?? |
cosmics |
L0DU (CaloCosmics) |
~10 Hz |
5(TAE) |
(1), (2) |
024076 (thurs.) |
x0 |
11k |
cosmics |
L0DU (CaloCosmics or Muon C) |
(15+2) Hz |
5(TAE) |
(3) (log_L0Trigger) |
024080 (thurs.) |
x0 |
67k |
cosmics |
L0DU (CaloCosmics or Muon C) |
(15+20) Hz |
5(TAE) |
(3) (log_Shift) , (log_L0Trigger) |
024090 (thurs.) |
x0 |
4k |
cosmics |
L0DU (CaloCosmics or Muon C) |
?? |
5(TAE) |
(3) (log_L0Trigger) |
024106 (thurs.) |
x0 |
|
cosmics |
L0DU (CaloCosmics or Muon C) |
?? |
5(TAE) |
(3) |
024109 (thurs.) |
x0 |
?? |
cosmics |
L0DU (CaloCosmics or Muon C) |
?? |
5(TAE) |
(3) |
024111 (thurs.) |
x0 |
?? |
cosmics |
L0DU (CaloCosmics or Muon C) |
?? |
5(TAE) |
A+C side for Ps/Spd |
024236 (thurs.) |
x0 |
?? |
cosmics |
L0DU (CaloCosmics) (4) |
?? |
5(TAE) |
PHYSICS_Cosmics recipe (log_Calo) |
024244 (thurs.) |
x0 |
?? |
cosmics |
L0DU (CaloCosmics) (4) |
?? |
5(TAE) |
PHYSICS_Cosmics recipe (log_Calo) |
(1) C-side of Ps/Spd without HV
(2) Hcal detector was moving
(3) With VELO C + IT + MUON C, besides full CALO + ODIN
(4) L0DU algorithm : S1 = (ECAL>20 and HCAL>5)
May 2008
Data (30/05/2008) [Ecal LED]
Data content :
- Ecal (A+C) LED
- each PMT flashing by 1/8 trigger rate
- default LEDTSB delay : 150ns
- nominal HV , nominal LED intensity, default FErecipe
- data available at : /daqarea/lhcb/data/2008/RAW/ECAL/PHYSICS_TimeScan/
- run coord. : Anatoli / Marie-Noelle
run |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
023630 |
1600/step |
LED scan |
calibration |
56Hz |
5(TAE) |
Ecal LED scan |
023633 |
1600/step |
LED scan |
calibration |
56Hz |
5(TAE) |
Ecal LED scan |
022635 |
20k |
LED |
calibration |
56 Hz |
5(TAE) |
Ecal LED |
022639 |
1600/step |
LED scan |
calibration |
56 Hz |
5(TAE) |
Ecal LED scan |
022642 |
20k |
LED |
calibration |
56 Hz |
5(TAE) |
Ecal LED |
022643 |
1600/step |
LED scan |
calibration |
56 Hz |
5(TAE) |
Ecal LED scan |
023650 |
1600/step |
LED scan |
calibration |
56Hz |
5(TAE) |
Ecal LED scan |
023691 |
20k |
LED |
calibration |
56Hz |
5(TAE) |
Ecal LED |
023694 |
1600/step |
LED scan |
calibration |
56Hz |
5(TAE) |
Ecal LED scan |
023699 |
20k |
LED |
calibration |
56Hz |
5(TAE) |
Ecal LED |
023700 |
1600/step |
LED scan |
calibration |
56Hz |
5(TAE) |
Ecal LED scan |
Data (29/05/2008) [Cosmic run]
Data content :
- Calo C-sides + Muon Q3 ODIN +L0DU
- run coord. : Patrick / Olivier C.
run |
mdf/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
023606 |
x0 |
?? |
cosmics |
L0DU (Muon only) |
few Hz |
5(TAE) |
Calo/Muon time aligned in DAQ (log_Shift) |
023604 |
x0 |
?? |
cosmics |
L0DU (Calo + Muon (1) ) |
7Hz(L0Calo) + 1Hz(L0Muon) |
5(TAE) |
L0Calo/L0Muon time aligned (log_Shift) ,(log_L0+Trigger) |
023626 |
x0 |
?? |
cosmics |
L0DU (L0hadron only (2) ) |
20Hz |
5(TAE) |
Calo A+C side nominal HV |
(1) L0 := [L0had(Et)>3 AND L0Elec(Et)>20] OR [L0mu(Pt)>0)
(2) L0 := [L0had(Et)>50 (~1GeV) Ecal+Hcal sum in TVB]
Data (27/05/2008) [Hcal LED scan]
- Hcal A+C-side
- Automatic FSM scan of LED timing
- 1600 evts/step
- each PMT flashing by 1/8 trigger rate
- data available at : /castor/cern.ch/grid/lhcb/data/2008/RAW/HCAL/PHYSICS_TimeScan
- run coord. : Anatoli / Marie-Noelle
run |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
023356 |
1600/step |
LED scan |
calibration |
56Hz |
5(TAE) |
HV nominal, delay 150ns, LED intensity nominal |
023375 |
1600/step |
LED scan |
calibration |
56Hz |
5(TAE) |
HV nominal, delay 150ns, LED intensity nominal |
023388 |
1600/step |
LED scan |
calibration |
56Hz |
5(TAE) |
HV nominal, delay 150ns, LED intensity nominal |
023404 |
1600/step |
LED scan |
calibration |
56Hz |
5(TAE) |
HV nominal, delay 150ns, LED intensity nominal |
023407 |
1600/step |
LED scan |
calibration |
56Hz |
5(TAE) |
HV nominal, delay 150ns, LED intensity nominal |
Data (23/05/2008) [Ecal LED scan]
- Ecal C-side
- Automatic FSM scan of LED timing
- 65 steps
- triggering channel by channel @ 56 Hz
- data available at : /daqarea/lhcb/data/2008/RAW/ECAL/PHYSICS|TimeScan/
- run coord. : Anatoli
run |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
023287 |
1k/step * 65 |
LED scan |
calibration |
56Hz |
5(TAE) |
HV nominal - LED intensity nominal - delay 150ns |
022319 |
100/step * 65 |
LED scan |
calibration |
56Hz |
5(TAE) |
HV nominal - LED intensity nominal - delay 150 ns |
Data (22/05/2008) [Cosmic run]
Data content :
- Calo A & C sides + ODIN +L0DU
- run coord. : Patrick/Olivier C.
Comments :
data are available on /daqarea and on the (new) castor path :
/castor/cern.ch/
grid /lhcb/data/2008/RAW/LHCb/PHYSICS/
run |
mdf/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
023230 |
x0 |
~14k |
cosmics |
L0DU (CaloCosmics) (3) |
few Hz |
3(TAE) |
(1) (log) |
023233 |
x0 |
~21k |
cosmics |
L0DU (CaloCosmics) (3) |
few Hz |
5(TAE) |
(2) |
023201 |
x0 |
~23k |
cosmics |
L0DU (CaloCosmics) (3) |
few Hz |
5(TAE) |
Ecal+Hcal+Prs (log) |
(1) Full detector
(2) Inner part of Prs/Spd disabled
TVB0 on A-side disabled
(3)
Warning : the L0DU algorithm has changed S1 = [had_Et > 7 && el_ET > 20] (was 3 for hadron)
Data (15&17/05/2008) [LED scan]
- Ecal C-side
- Automatic FSM scan of LED timing
- zigzag 1/8 frequency
- data available at : /daqarea/lhcb/data/2008/RAW/ECAL/PHYSICS|TimeScan/
- run coord. : Anatoli
run |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
023035 |
800/step * 13 |
LED scan |
calibration |
56Hz |
5(TAE) |
HV nominal - LED intensity nominal - delay 150ns |
023036 |
400/step * 250 |
LED scan |
calibration |
56Hz |
5(TAE) |
HV nominal - LED intensity nominal - delay 150 ns |
022984 |
400/step * 200 |
LED scan |
calibration |
56Hz |
5(TAE) |
HV nominal - LED intensity 2000cnt- delay 150 ns |
Data (14/05/2008) [LED scan]
- E/Hcal both sides
- Automatic FSM scan of LED timing
- 10k event per step
- run coord : Anatoli
run |
mdf/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
022890 |
x0 |
10k/step |
LED scan |
calibration |
?? |
5(TAE) |
Ecal scan |
022903 |
x0 |
10k/step |
LED scan |
calibration |
?? |
5(TAE) |
Hcal scan |
Data (06/05/2008) [Global commissioning cosmics run]
Data content :
- Ecal/Hcal/Prs A-side + ODIN
- run coord : Olivier C.
run |
mdf/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
022584 |
x0 |
?? |
cosmics |
L0DU (CaloCosmics) |
few Hz |
13(TAE) |
-- |
Data (06/05/2008) [Global commissioning cosmics run]
Data content :
- Ecal/Hcal/Prs A-side + ODIN
run |
mdf/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
022584 |
x0 |
?? |
cosmics |
L0DU (CaloCosmics) |
few Hz |
13(TAE) |
-- |
022579 |
x0 |
?? |
cosmics |
L0DU (CaloCosmics) |
few Hz |
13(TAE) |
-- |
022578 |
x0 |
?? |
cosmics |
L0DU (CaloCosmics) |
few Hz |
13(TAE) |
-- |
022574 |
x0 |
?? |
cosmics |
L0DU (CaloCosmics) |
few Hz |
13(TAE) |
-- |
022547 |
x0 |
?? |
cosmics |
L0DU (CaloCosmics) |
few Hz |
13(TAE) |
-- |
022530 |
x0 |
?? |
cosmics |
L0DU (CaloCosmics) |
few Hz |
5(TAE) |
(3) |
022476 |
x0 |
?? |
cosmics |
L0DU (CaloCosmics) |
few Hz |
5(TAE) |
(2) |
022465 |
x0 |
?? |
cosmics |
L0DU (CaloCosmics) |
few Hz |
5(TAE) |
(1) |
(1) Trigger rate exploded after few kEvent have been stored (due to Ecal channel becoming hot) - Only first few kEvent are cosmics.
(2) same configuration than run 22465 but masking hot calo channel
(3) LED interspersed
Data (02/05/2008) [FSM automatic scan : Ecal A-side]
Data content :
- Ecal A-side + ODIN
-> the crate 17 is back (work after hard reset)
Comments
- Automatic FSM scan of LED timing : 12 steps of 2 ns;1200 ev/step, LED flashing freq. 1/16 of trigger rate
- Data are available in the online system at : /daqarea/lhcb/data/2008/RAW/HCAL/PHYSICS|TimeScan/[RUNID]
- also available on CASTOR at : /castor/cern.ch/user/k/konoplia/
-
Warning : the scan step is not available in the ODIN data.
run |
mdf/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
022212 |
x0 |
12x1200 |
LED |
calibration |
56Hz |
5(TAE) |
HV normal, LEDTSB_del=150 ns; LED_intens = 1.00*normal cnt |
022215 |
x0 |
12x1200 |
LED |
calibration |
56Hz |
5(TAE) |
HV normal, LEDTSB_del=150 ns; LED_intens = 0.95*normal cnt |
022217 |
x0 |
12x1200 |
LED |
calibration |
56Hz |
5(TAE) |
HV normal, LEDTSB_del=150 ns; LED_intens = 0.80*normal cnt |
Data (01/05/2008) [FSM automatic scan : Hcal/Ecal A-side]
Data content :
- Hcal A-side + ODIN
Comment
- Automatic FSM scan of LED timing : 12 steps of 2 ns; 800 ev/step; LED flashing freq. 1/8 of trigger rate
- Data are available in the online system at : /daqarea/lhcb/data/2008/RAW/HCAL/PHYSICS|TimeScan/[RUNID]
- also available on CASTOR at : /castor/cern.ch/user/k/konoplia/
-
Warning : the scan step is not available in the ODIN data. This needs to be fixed with Richard.
run |
mdf/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
022160 |
x0 |
12x800 |
LED |
calibration |
56Hz |
5(TAE) |
HV normal, LEDTSB_mean_delay=150 ns; LED_intens=1800 cnt |
Data content :
- Ecal A-side + ODIN
-> the crate 17 was excluded, due to a CROC problem
-> the EC19FEB14 board was excluded, due to it was unreachable by SPECS
- Automatic FSM scan of LED timing : 12 steps of 2 ns;1200 ev/step, LED flashing freq. 1/16 of trigger rate
- Data are available in the online system at : /daqarea/lhcb/data/2008/RAW/HCAL/PHYSICS|TimeScan/[RUNID]
- also available on CASTOR at : /castor/cern.ch/user/k/konoplia/
-
Warning : the scan step is not available in the ODIN data. This needs to be fixed with Richard.
run |
mdf/root |
#sequences |
data |
Trigger type |
rate |
BX sequence |
comment |
022166 |
x0 |
12x1200 |
LED |
calibration |
56Hz |
5(TAE) |
HV normal, LEDTSB_del=150 ns; LED_intens=normal cnt |
022167 |
x0 |
12x1200 |
LED |
calibration |
56Hz |
5(TAE) |
HV normal, LEDTSB_del=170 ns; LED_intens=normal cnt |
022168 |
x0 |
12x1200 |
LED |
calibration |
56Hz |
5(TAE) |
HV normal, LEDTSB_del=130 ns; LED_intens=normal cnt |
April 2008
Data (10/04/2008) [FSM automatic scan]
Data content :
- Hcal C-side + ODIN
Comment
- Automatic FSM scan of LED timing : 12 steps with 2 ns STEP, 400 and 800 events per step.
run |
mdf/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
comment |
21559 |
x0 |
12x800 |
LED |
calibration |
? |
5(TAE) |
HV=1300 V (2500 cnt) LED ~1500 cnt |
21560 |
x0 |
12x800 |
LED |
calibration |
? |
5(TAE) |
HV=1400 V (2700 cnt) LED ~1500 cnt |
21561 |
x0 |
12x800 |
LED |
calibration |
? |
5(TAE) |
HV=1050 V (2000 cnt) LED ~1500 cnt |
21562 |
x0 |
12x800 |
LED |
calibration |
? |
5(TAE) |
HV for Gain 200 000, LED ~1500 cnt. |
21563 |
x0 |
12x800 |
LED |
calibration |
? |
5(TAE) |
HV for Gain 200 000, LED ~1500*0.9 cnt. |
21564 |
x0 |
12x800 |
LED |
calibration |
? |
5(TAE) |
HV for Gain 200 000, LED ~1500*0.81 cnt. |
21565 |
x0 |
12x800 |
LED |
calibration |
? |
5(TAE) |
HV for Gain 200 000, LED ~1500*0.81 cnt. (PHYSICS recipe for TELL1) |
Data (04/04/2008) [Global commissioning session]
Data content :
- Hcal/Ecal/(Prs/Spd) C-side + ODIN + L0DU + L0Calo (+ L0Muon + Muon)
Comment
- ecTell02 disabled (configuration problem)
WARNING : new PRS TELL1 firmware -> patch for Prs/Spd L0-bit decoding NO LONGER NEEDED
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
021325_0000002345.raw |
x000 |
?? |
cosmics |
L0 (Calo) |
4 Hz |
7(TAE) |
100% |
(1) |
021326_0000002346.raw |
x000 |
?? |
cosmics |
L0 (Calo) |
4 Hz |
7(TAE) |
100% |
(1) |
021327_0000002348.raw |
x000 |
?? |
cosmics |
L0 (Calo) |
4 Hz |
7(TAE) |
100% |
(1) |
021328_0000002349.raw |
x000 |
?? |
cosmics |
L0 (Calo) |
4 Hz |
7(TAE) |
100% |
(1) |
021329_0000002350.raw |
x000 |
?? |
cosmics |
L0 (Calo) |
4 Hz |
5(TAE) |
100% |
(1) |
021330_0000002351.raw |
x000 |
?? |
cosmics |
L0 (Calo) |
4 Hz |
5(TAE) |
100% |
(1) |
021351_0000002377.raw |
x000 |
?? |
cosmics |
L0 (Calo) |
4 Hz |
5(TAE) |
100% |
(1) |
021354_0000002380.raw |
x000 |
?? |
cosmics |
L0 (Calo) |
4 Hz |
5(TAE) |
100% |
(1) |
021355_0000002381.raw |
x000 |
?? |
cosmics |
L0 (Calo) |
4 Hz |
5(TAE) |
100% |
(1) |
021390_0000002426.raw |
x000 |
?? |
cosmics |
L0 (Calo) |
4 Hz |
13(TAE) |
100% |
(1) |
021391_0000002427.raw |
x000 |
?? |
cosmics |
L0 (Calo) |
4 Hz |
13(TAE) |
100% |
(1) |
021394_0000002433.raw |
x000 |
?? |
cosmics |
L0 (Calo or Muon) |
5 Hz |
13(TAE) |
100% |
(2) |
021395_0000002434.raw |
x000 |
?? |
cosmics |
L0 (Calo or Muon) |
5 Hz |
13(TAE) |
100% |
(2) |
021396_0000002435.raw |
x000 |
?? |
cosmics |
L0 (Muon) |
1 Hz |
13(TAE) |
100% |
(3) |
(1) Standard L0 CaloCosmics trigger ( S1=(electron>20) && (hadron>3) )
(2) Add L0 MuonCosmics trigger (S2 = (Muon1 >0) ) with L0MuonCandidate from Q3
Muon activity seen in Prev1 (1BX shift)
Relative trigger rate : Calo + Muon ~ 4 + 1 Hz
(3) L0 MuonCosmics trigger alone
Rate ~ 1Hz - cosmics purity expected ~ 10%
Correlated cosmics activities seen in Ecal/Hcal for ~< 1/10 trigger
March 2008
Data (03/19/2008) [Hcal C-side LED scan]
Data content :
- Hcal (C-side) + ODIN
Comment
- Hcal C-side Led scan with different configuration of gain or led light
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
hcaltest_1903_0006 |
xx00 |
~2k |
LED |
calibration |
115 Hz |
5(TAE) |
100% |
Nominal setting |
hcaltest_1903_0007 |
xx00 |
~2k |
LED |
calibration |
115 Hz |
5(TAE) |
100% |
mod#14 Gain=90% |
hcaltest_1903_0008 |
xx00 |
~2k |
LED |
calibration |
330 Hz |
5(TAE) |
100% |
mod#14 Gain=95% |
hcaltest_1903_0009 |
xx00 |
~2k |
LED |
calibration |
250 Hz |
5(TAE) |
100% |
mod#14 LED=~80% |
Data (03/14/2008) [Ecal/Hcal A-side Cosmics in Global Commissioning session]
Description : Calo Cosmics run during global commissioning session
Data content :
- Ecal/Hcal (A-side) + L0Calo + L0DU + ODIN
Comment
- ecTell14 was disabled (problem : seems to not receive TTC)
- L0Calo banks are not OK --> Don't use them (intermediate firmware)
The data files are available on the online system at :
/daqarea/lhcb/data/2008/RAW/LHCb/PHYSICS/20876/
and transfered on /castor/cern.ch/lhcb/online/point8/2008/RAW/LHCb/PHYSICS/20876/
Duplicate is also available at /castor/cern.ch/user/o/odescham/data/commissioning/com0308/
WARNING THESE ARE MDF DATA
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
020876_0000001440.raw |
x000 |
15000 |
cosmics |
L0 |
~5 Hz |
3(TAE) |
100% |
(log) |
Data (03/12/2008)
Ecal LED scan
Description : Ecal LED scan
HV = nominal
Data content :
- Ecal C-side
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
ecal_ledscan_1203-0001 |
xx0x |
~5k |
LED |
calibration |
110 Hz |
5(TAE) |
100% |
Delay=130 |
ecal_ledscan_1203-0002 |
xx0x |
~3k |
LED |
calibration |
110 Hz |
5(TAE) |
100% |
Delay=140 |
ecal_ledscan_1203-0003 |
xx0x |
~3k |
LED |
calibration |
110 Hz |
5(TAE) |
100% |
Delay=150 |
ecal_ledscan_1203-0004 |
xx0x |
~3k |
LED |
calibration |
110 Hz |
5(TAE) |
100% |
Delay=160 |
ecal_ledscan_1203-0005 |
xx0x |
~3k |
LED |
calibration |
110 Hz |
5(TAE) |
100% |
Delay=170 |
Data (03/11/2008)
Ecal LED scan
Description : Ecal LED scan
HV = nominal
Data content :
- Ecal C-side
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
ecal_ledscan_1103-0001 |
xx0x |
3.4k |
LED |
calibration |
36 Hz |
5(TAE) |
100% |
Delay=130 |
ecal_ledscan_1103-0002 |
xx0x |
~3k |
LED |
calibration |
36 Hz |
5(TAE) |
100% |
Delay=140 |
ecal_ledscan_1103-0003 |
xx0x |
~3k |
LED |
calibration |
112 Hz |
3(TAE) |
100% |
Delay=140 |
ecal_ledscan_1103-0004 |
xx0x |
~3k |
LED |
calibration |
36 Hz |
5(TAE) |
100% |
Delay=150 |
ecal_ledscan_1103-0005 |
xx0x |
~3k |
LED |
calibration |
56 Hz |
5(TAE) |
100% |
Delay=160 |
ecal_ledscan_1103-0006 |
xx0x |
~10k |
LED |
calibration |
56 Hz |
5(TAE) |
100% |
Delay=170 |
Test transmission errors
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
ecal_ledscan_1103-0007 |
xx0x |
49k |
LED |
calibration |
280 Hz |
5(TAE) |
100% |
MEPf=5 (1) |
ecal_ledscan_1103-0008 |
xx0x |
78k |
LED |
calibration |
380 Hz |
3(TAE) |
100% |
MEPf=3 (1) |
ecal_ledscan_1103-0009 |
xx0x |
54k |
LED |
calibration |
160->400 Hz |
7(TAE) |
100% |
MEPf=3 (2) |
ecal_ledscan_1103-0010 |
xx0x |
50k |
LED |
calibration |
380 Hz |
1 |
100% |
MEPf=5 (1) |
ecal_ledscan_1103-0011 |
xx0x |
50k |
LED |
calibration |
380 Hz |
1 |
100% |
MEPf=3 (1) |
ecal_ledscan_1103-0012 |
xx0x |
55k |
LED |
calibration |
380 Hz |
3(TAE) |
100% |
MEPf=3 (2) |
(1) Bad header observed in decoding data
(2) NO ERROR IN DECODING
Prs noise
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
prs_1103-0001 |
xx0x |
??k |
noise |
calibration |
140 Hz |
5(TAE) |
100% |
Prs0 only |
prs_1103-0002 |
xx0x |
??k |
noise |
calibration |
140 Hz |
5(TAE) |
100% |
Prs0 + Prs2 |
Data (03/07/2008) [Ecal/Hcal/Prs Cosmics in Global Commissioning session]
Description : Calo Cosmics run during global commissioning session
Data content :
- more than 100 TELL1 among which Ecal/Hcal/Prs/L0Calo/L0DU/ODIN
data file available on the online system at :
/daqarea/lhcb/data/2008/RAW/LHCb/PHYSICS/20XXX/
and transfered on /castor/cern.ch/lhcb/online/point8/2008/RAW/LHCb/PHYSICS/20XXX/
(with XXX= 769, 772, 774 & 775)
WARNING THESE ARE MDF DATA
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
020769_0000001364.raw |
x000 |
? |
cosmics |
L0 |
~5x4 Hz |
5(TAE) |
100% |
Prs VFE Off |
020772_0000001372.raw |
x000 |
? |
cosmics |
L0 |
~5x4 Hz |
5(TAE) |
100% |
Prs VFE On |
020774_0000001376.raw |
x000 |
? |
cosmics |
L0 |
~5x4 Hz |
5(TAE) |
100% |
Prs VFE On (1) |
020775_0000001377.raw |
x000 |
? |
cosmics |
L0 |
~5x4 Hz |
5(TAE) |
100% |
Prs VFE On |
New : the TAE mode does work with mdf data format
(1) : Warning : problem with TAE sequence shifted by 1 BX from event 3626 (single trigger event inserted in a sequence).
For TAE analysis only use first 3600 events.
Data (03/06/2008) [Ecal/Hcal/Prs Cosmics in Global Commissioning session]
COSMICS EVENTS OBSERVED IN COINCIDENCE IN HCAL,ECAL and Prs for the first time
Description : Calo Cosmics run during global commissioning session
Data content :
- Ecal/Hcal/Prs/L0Calo/L0DU/ODIN
Prs latency 137
data file available on the online system at :
/daqarea/lhcb/data/2008/RAW/LHCb/PHYSICS/20716/
and transfered on /castor/cern.ch/lhcb/online/point8/2008/RAW/LHCb/PHYSICS/20716/
WARNING THESE ARE MDF DATA
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
020716_0000001248.raw |
x000 |
5613 |
cosmics |
L0 |
~4 Hz |
3(TAE) |
100% |
|
Warning : the TAE mode does not work with global commissioning system (identified ODIN problem) -> only the T0 sample is stored in tha avobe file
Data (03/05/2008) [Cosmics run]
Description : Calo Cosmics run
Data content :
- Ecal/Hcal/Prs/L0Calo/L0DU/ODIN
Ecal/Hcal trigger on outer C-side
Prs timing scan to try to see cosmics in coincidence
Actual pedestal subtraction on Prs0 - global -200 subtration on other Prs
Data files are available on /castor/cern.ch/lhcb/online/point8/calo/com0208
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
calo_cosmics_0503_0001 |
x000 |
few k |
cosmics |
L0 |
~6 Hz |
5(TAE) |
100% |
Prs latency 144 |
calo_cosmics_0503_0002 |
x000 |
few k |
cosmics |
L0 |
~6 Hz |
5(TAE) |
100% |
Prs latency 146 |
calo_cosmics_0503_0003 |
x000 |
few k |
cosmics |
L0 |
~6 Hz |
5(TAE) |
100% |
Prs latency 142 |
calo_cosmics_0503_0004 |
x000 |
few k |
cosmics |
L0 |
~6 Hz |
5(TAE) |
100% |
Prs latency 135 |
calo_cosmics_0503_0005 |
x000 |
few k |
cosmics |
L0 |
~6 Hz |
5(TAE) |
100% |
prs latency 139 |
Data (03/03/2008) [Cosmics Global Commissioning ]
Description : Calo Cosmics run during global commissioning session
Data content :
- Ecal/Hcal/L0Calo/L0DU/ODIN
Two mdf data files stored on /daqarea/lhcb/data/2008/RAW/LHCb/PHYSICS/20664/*raw
and /daqarea/lhcb/data/2008/RAW/LHCb/PHYSICS/20664/*raw
on the Online system.
Files also transfered on /castor/cern.ch/lhcb/online/point8/calo/com0308/
WARNING THESE ARE MDF DATA
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
020664_0000000938.raw |
x000 |
fewk |
cosmics |
L0 |
~10 Hz |
3(TAE) |
100% |
|
020665_0000000942.raw |
x000 |
fewk |
cocmics |
L0 |
~10 Hz |
3(TAE) |
100% |
|
January/February 2008
Data (02/21/2008) [Prs C-side commissioning ]
Description : Prs C-side
Data content :
- Prs crates : Prs0 -> Prs3
- ODIN-FE
- Last file has been taken w/ +/- 1 event around the current BX. All PRS crates were in.
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
prs_2102_0001 |
xxx0 |
1.4k |
noise |
periodic |
10 Hz |
1 |
100% |
Prs1 |
prs_2102_0002 |
xxx0 |
1.1k |
noise |
periodic |
10 Hz |
1 |
100% |
Prs3 |
prs_2102_0003 |
xxx0 |
2.k |
noise |
periodic |
10 Hz |
1 |
100% |
Prs0,1,3 |
prs_2202_0001b |
xxx0 |
?? |
noise |
periodic |
10 Hz |
3 (TAE) |
100% |
Prs0,1,2,3 |
Data (02/13/2008) [E/HCAL LED SCAN + A-side commissioning + COSMICS RUN ]
E/HCal LED scan
Description : ECAL / HCAL LED scan
Data content :
- Ecal crates : 13 Tell1 banks: Full C-side (first time complete)
- Hcal crates : 4 Tell1 banks : Full C-side (first time complete)
- ODIN-FE : 1 bank
--> Total of 18 banks
Setup :
TFC = ODIN10
NEW Tell1 f/w : 2.39
MEP factor = 5
TAE window = +- 2
Comments :
- Still problem with ecTELL07 data (time to time wrong FEBId/CrateId in FEB#5 header ~1% of events
- swapping ecTELL07 and ecTELL08 input from CROC showed the problem is related to the ecTELL07 not the input data
HV setting :
Ecal HV : nominal
Hcal HV : 1.3kV
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
calo_1302_0001 |
xx0x |
7.0k |
LED |
calibration |
50 Hz |
5 (TAE) |
100% |
Ecal : 130ns Hcal :150ns |
calo_1302_0002 |
xx0x |
3.6k |
LED |
calibration |
50 Hz |
5 (TAE) |
100% |
Ecal : 130ns Hcal :150ns |
calo_1302_0003 |
xx0x |
2.1k |
LED |
calibration |
50 Hz |
5 (TAE) |
100% |
Ecal : 140ns Hcal :152ns |
calo_1302_0004 |
xx0x |
3.2k |
LED |
calibration |
50 Hz |
5 (TAE) |
100% |
Ecal : 150ns Hcal :154ns |
calo_1302_0005 |
xx0x |
2.0k |
LED |
calibration |
50 Hz |
5 (TAE) |
100% |
Ecal : 160ns Hcal :156ns |
calo_1302_0006 |
xx0x |
2.4k |
LED |
calibration |
50 Hz |
5 (TAE) |
100% |
Ecal : 170ns Hcal :158ns |
Ecal HV : 630V
Hcal HV : 1.3kV
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
calo_1302_0007 |
xx0x |
3.0k |
LED |
calibration |
50 Hz |
5 (TAE) |
100% |
Ecal : 170ns Hcal :160ns |
calo_1302_0008 |
xx0x |
2.2k |
LED |
calibration |
50 Hz |
5 (TAE) |
100% |
Ecal : 160ns Hcal :162ns |
calo_1302_0009 |
xx0x |
3.7k |
LED |
calibration |
50 Hz |
5 (TAE) |
100% |
Ecal : 150ns Hcal :164ns |
calo_1302_0010 |
xx0x |
6.3k |
LED |
calibration |
50 Hz |
5 (TAE) |
100% |
Ecal : 150ns Hcal :164ns |
calo_1302_0011 |
xx0x |
2.0k |
LED |
calibration |
50 Hz |
5 (TAE) |
100% |
Ecal : 140ns Hcal :166ns |
calo_1302_0012 |
xx0x |
3.0k |
LED |
calibration |
50 Hz |
5 (TAE) |
100% |
Ecal : 130ns Hcal :168ns |
Noise DAQ including Ecal-A
Description : ECAL / HCAL
Data content :
- Ecal crates : 26 TELL1 banks: Full ECAL (except fist file : C-side only without ecTELL07)
- Hcal crates : 4 Tell1 banks : Full C-side
- ODIN-FE : 1 bank
--> Total of 31 banks
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
calo_1302_0013 |
xx00 |
?? |
noise |
periodic |
50 Hz |
1 |
100% |
C-side only - ecTELL07 disabled |
fullcalo_1302_0017 |
xx00 |
?? |
noise |
periodic |
50 Hz |
1 |
100% |
C-side + Ecal-A (1) |
fullcalo_1302_0018 |
xx00 |
?? |
noise |
periodic |
50 Hz |
1 |
100% |
C-side + Ecal-A (2) |
(1) mis-cabling detected on A-side : ecTELL14 and ecTELL21 (partially swapped)
(2) software 'correction' for miscabling : swap ecTELL14 and ecTELL21 input from CROC. swap configuration. Still problems
Cosmics run
Description : ECAL / HCAL (/ Prs) cosmics run
Data content :
- Ecal crates : 13 Tell1 banks: Full C-side
- Hcal crates : 4 Tell1 banks : Full C-side
- L0DU: 1 Tell1 bank
- L0Calo: 2 Tell1 banks
- ODIN-FE : 1 bank
--> Total of 21 banks
Setup :
TFC = ODIN10
NEW Tell1 f/w : 2.39
MEP factor = 5
HV setting ("cosmics setting") :
Ecal HV : Gain = 3.10^5
Hcal HV : Gain = 2.10^5
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
calo_cosmics_1302_0018 |
xx00 |
?? |
cosmics |
L0-external (1) |
1kHz |
1 |
100% |
(2) |
calo_cosmics_1302_0019 |
xx0x |
20k |
cosmics |
L0-external (1) |
10-15Hz |
1 |
100% |
(3) High Quality |
calo_cosmics_1302_0024 |
xx0x |
?? |
cosmics |
L0-external (1) |
10-15 Hz |
5 (TAE) |
100% |
(4) |
(1) Default L0DU algorithm (
(see below))
(2) Data run used to detect hot channels (Ecal crate 14 FEB#7,8,9) - L0Calo calibration constant = 128
(3) hot channels removed from L0Calo path. L0Calo calibration constant down to 127.
LARGE FRACTION OF CLEAR COSMICS EVENTS (~4/5) @ 10-15 Hz
(4) Attempt to get cosmics in TAE mode. Although ODIN bank have BCID that indicate that there is shift in TAE sequence, the data should be OK (
L0DU bank provide correct BCID while ODIN DAQ seems to be wrong)
Data (02/07 & 08/2008) [COSMICS RUN]
Description : ECAL / HCAL /L0Calo/L0DU commissioning
Data content :
- Ecal crates : 12 Tell1 banks: Full C-side except ecTell12 (part of crate 13 & 14)
- Hcal crates : 3 Tell1 banks : Full C-side except hcTell04 (half of crate 24)
- ODIN-FE : 1 bank
- L0DU: 1 Tell1 banks
- L0Calo : 2 Tell1 banks
--> Total of 18 banks
Setup :
NEW Tell1 f/w : 2.39
Ecal HV = cosmics nominal (Constant gain)
Hcal HV = cosmics nominal (Constant gain)
Comments :
- Some hot channels (responsible of high trigger rate) were identified and removed from the trigger path :
-- Ecal crate#14 FEB#7 channel 19
-- Ecal crate#14 FEB#8 channel 29,30 & 31
When these channels are removed from the trigger the 'cosmics' trigger rate decrease from ~220 Hz to ~20Hz.
Seems to remains few (less) hot channels in Ecal and Hcal. To be identified with detailed study of the output files.
- ADC=-256 in PIN-FEB (Ecal crate 11) have disappeared but observed intermittent ADC=-256 problem in Ecal crate#10 FEB#8 (very few events - occurs on several channels)
- L0ADC.vs.ADC have now a much correct shape although the configuration is the same as for 06/07 february run. Seems to remain some problems in crate 11.
- L0DU algorithm as for 06/07 february run
(see below).
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
calo_cosmics_0702_0001 |
xx0x |
300 |
cosmics |
L0 |
20 Hz |
1 |
5% |
(1) |
calo_cosmics_0702_0002 |
xx0x |
370 |
cosmics |
L0 |
20 Hz |
1 |
5% |
(1) |
calo_cosmics_0702_0003 |
xx0x |
370 |
cosmics |
L0 |
20 Hz |
1 |
5% |
(2) |
calo_cosmics_0702_0004 |
xx0x |
?? |
cosmics |
L0 |
20 Hz |
1 |
5% |
(2) |
calo_cosmics_0702_0005 |
xx0x |
?? |
cosmics |
L0 |
20 Hz |
5 (TAE) |
5% |
(3) |
calo_cosmics_0702_0006 |
xx0x |
?? |
cosmics |
L0 |
20 Hz |
5 (TAE) |
100% |
(3) |
(1) - MEP factor = 3 - No Timing Trigger Bit request from L0DU
(2) - MEP factor = 5 with attempt to use
L0DU Timing Trigger but bad ODIN configuration (Timing window =0 !).
So should be identical to (1).
All TELL1 have been reset betweeen run 0702_003 and 0702_004
(3) - Mep factor = 5 - L0DU Timing Trigger activated with Window = +-2.
Problem with Timing mode seen on last file : 3 single BX event are inserted in a MEP and the last BX of the sequence is lost
--> MEP and BX sequence does not match after this problem occurs
Data (02/06 & 07/2008) [COSMICS RUN]
Description : ECAL / HCAL / L0DU commissioning
Data content :
- Ecal crates : Full C-side except ecTell12 (part of crate 13 & 14)
- Hcal crates : Full C-side except hcTell04 (half of crate 24)
- ODIN-FE
- L0DU
Setup :
NEW Tell1 f/w : 2.39
MEP factor = 3
Ecal HV = nominal
Hcal HV = 1.3 kV
Comments :
The previous problem ADC=-256 for Ecal FIXED for most of CROC
was due to bad default configuration of Ecal FEBs -> Frederic should fix the default recipes.
- the problem remains for pin-DIODE FEB in Ecal crate 11.
- In addition : intermittent problem with FEB/Crate code = 0 in header for crate 11 FEB6
- In addition : L0ADC vs ADC is not correct for some data from crate 11.
All these problems are probably related to a bad synchro setting for CROC/FEB in crate 11
Again difficulties to configure the Croc/FEB via PVSS
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
calo_cosmics_0602_0001 |
xxxx |
?? |
cosmics |
L0 (1) |
30 Hz |
1 |
100% |
(2) |
calo_cosmics_0602_0002 |
xxxx |
?? |
cosmics |
L0 (1) |
< 1 Hz |
1 |
100% |
(3) |
calo_cosmics_0602_0003 |
xxxx |
?? |
cosmics |
L0 (1) |
< 1 Hz |
1 |
100% |
(3) |
calo_periodic_0602_0002 |
xx0x |
?? |
noise |
periodic |
100Hz |
1 |
5% |
(3) |
calo_calib_0602_0002 |
xxxx |
?? |
LED |
calibration |
100Hz |
1 |
15% |
(3) |
(1) External L0 : C1=ET(electron)>20 and C2=ET(had)>3 [same as for cosmics run on 12/20/2007]
(2) : Pb ADC = -256 in Ecal full FEBs or few channels/FEB. Hot channels (inner Ecal) leads to large trigger rates.
(3) : ADC pb fixed (except for PIN-FEB in Ecal crate 11). Inner Ecal removed from trigger path.
Root file (with ntuple and views) are available at /castor/cern.ch/lhcb/online/point8/calo/ntp0208/
Hbook files are available at : /castor/cern.ch/user/o/odescham/data/commissioning/ntp0208
-----
Description : PRS0 / L0DU standalone commissioning
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
prs_0602_0004 |
xxxx |
330 |
noise |
periodic |
112Hz |
1 |
5% |
Prs0 + ODIN |
l0du_0602_0003 |
xx00 |
?? |
noise |
periodic |
112Hz |
1 |
5% |
L0DU + ODIN |
Data (02/04 & 05/2008)
Description : ECAL and/or HCAL
Data content :
- Ecal crates : Full C-side except ecTell12 (part of crate 13 & 14)
- Hcal crates : Full C-side except hcTell04 (half of crate 24)
- ODIN-FE
NEW Tell1 f/w : 2.39
MEP factor = 3
EHcal HV = nominal LED ON
Comments :
latest TELL1 f/w is OK : no longer need of patch for decoding. no longer problem with L0ADC/ADC alignment.
problem with Ecal :: ADC=-256 on all channels for 4 FEB and on some channels of 1O other FEB.
Difficulties to configure the Croc/FEB via PVSS
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
ecal_0402_0001 |
xx00 |
118k |
noise |
periodic |
112Hz |
1 |
1% |
Ecal alone |
ecal_0402_0002 |
xx00 |
?? |
LED |
calibration |
68Hz |
3 (TAE) |
1% |
Ecal alone |
ecal_0502_0001 |
xx00 |
?? |
LED |
calibration |
112Hz |
3 (TAE) |
5% |
Ecal alone |
calo_0502_0002 |
xx00 |
?? |
LED |
calibration |
56Hz |
3 (TAE) |
5% |
Ecal + Hcal |
calo_0502_0003 |
xx00 |
?? |
LED |
calibration |
56Hz |
3 (TAE) |
5% |
Ecal + Hcal |
calo_0502_0004 |
xx00 |
?? |
LED |
calibration |
56Hz |
3 (TAE) |
5% |
Hcal alone |
calo_0502_0005 |
xx00 |
?? |
LED |
calibration |
300Hz |
3 (TAE) |
5% |
Hcal alone |
Description : Prs
Data content :
- Prs0
- ODIN-FE
Comments :
problem with FEB6 (known) and FEB1 (new !)
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
prs_0502_0001 |
xx00 |
?? |
noise |
periodic |
300Hz |
1 |
5% |
Prs |
Data (01/31/2008)
Description : PRS C-side commissioning
Data content :
- Prs0 (except FEB6 that has problem with crateID) -> psTell01
- ODIN-FE
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
prs_3101_0001 |
xx00 |
?? |
noise |
calibration |
112Hz |
3(TAE) |
1% |
Prs alone |
December 2007
Data (12/20/2007) [COSMICS RUN]
Description : ECAL & HCAL & L0Calo C-side commissioning
Data content :
Ecal crates : 9 + 10 --> TELL1 : 03 - 04 - 05 -06
Hcal crates : 22 --> TELL1 : 01 - 02
TrgCalo : TELL1 02
ODIN-FE
Comments :
Temperature problem with Ecal crate 8 solved :: Power Supply replaced but problem with decoding of TELL1 -> tell01 & tell02 are removed
problem with Ecal crate 12 :: still cannot switch on although the supply has been replaced (control problem ?: the power supply was not
plugged correctly)
problem with Ecal Tell12 :: still cannot be configured
Configuration of FE Crates done with CAT (PVSS not adapted to new CROC firmware)
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
calo-l0_1220_0003 |
mdf |
1 |
PMT Ecal (Gain 3.10^5)/Hcal (1.3kV) ON |
External L0 : ET(electron)>20 |
48 Hz (downscaling of 1% in L0DU) |
1 |
100% |
|
calo-l0_1220_0004 |
mdf |
1 |
PMT Ecal (Gain 3.10^5)/Hcal (1.3kV) ON |
External L0 : ET(electron)>20 |
48 Hz (downscaling of 1% in L0DU) |
1 |
100% |
|
calo-l0_1220_0010 |
mdf |
1 |
PMT Ecal (Gain 3.10^5)/Hcal (1.3kV) ON |
External L0 : ET(electron)>20 and ET(had)>3 |
4 Hz |
1 |
100% |
|
calo-l0_1220_0011 |
mdf |
1 |
PMT Ecal (Gain 3.10^5)/Hcal (1.3kV) ON |
External L0 : ET(electron)>20 and ET(had)>3 |
4 Hz |
1 |
100% |
|
calo-l0_1220_0020 |
mdf |
5 |
PMT Ecal (Gain 3.10^5)/Hcal (1.3kV) ON |
External L0 : ET(electron)>20 and ET(had)>3 |
4 Hz |
1 |
100% |
|
Data (12/19/2007)
Description : ECAL & HCAL & L0Calo C-side commissioning
Data content :
Ecal crates : 9 + 10 + 11 +13 +14 --> TELL1 : 03 - 04 - 05 -06 - 07 - 08 - 11 - 13
Hcal crates : 22 & 23 --> TELL1 : 01 - 02 - 03
TrgCalo : TELL1 01 - 02
ODIN-FE
(L0DU was first added in the banks list but cannot support the consecutive BX mode : need an f/w upgrade)
Comments :
problem with Ecal crate 8 :: too hot (ecTell01 & ecTell02)
problem with Ecal crate 12 :: cannot switch on ( ecTell09 & ecTell10)
problem with Ecal Tell12 :: cannot be configured
problem with Hcal Tell04 :: cannot be configured
Hcal : HV = 1.3kV LED ON delay 140 ns
Ecal : HV = nominal No LED delay 140 ns
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
calo-l0_1219_0001 |
xx00 |
??? |
LED HCAL |
calibration |
560Hz |
5(TAE) |
100% |
Hcal LED ON |
Data content : same as above with PRS0 added
Comment : problem Prs data cannot be decoded correctly : seems to be due to TELL1 firmware version 2.1.
Test with old firmware (2.0) : better decoding except for FEB6 that need to be desactivated
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
calo-l0_1219_0002 |
xx00 |
??? |
noise |
periodic |
10Hz |
5(TAE) |
100% |
Tell1 f/w v2.1 |
Data content : Ecal crates 9 (ecTell03 & ecTell04) + TrgCalo (Tell 01 & 02) + ODIN-FE
Comment : problem with FE-board Crate9#4 & Crate9#6 that produce noise around ADC=512. Need to be restarted
Trigger : L0(photon) > 50
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
ecal-l0_1219_0002 |
xxxx |
~2000 |
no LED |
external L0 - L0(g)>50 |
~1Hz |
1 |
100% |
pb : ADC = -256 on channel 31 |
ecal-l0_1219_0003 |
xx00 |
??? |
no LED |
external L0 - L0(g)>50 |
~1Hz |
1 |
100% |
last channel pb fixed |
Data (12/12/2007)
Description : Hcal C-side commissioning
Data content : LED data - Full Hcal C-side (4 TELL1 - crate 22 & 23) & ODIN-FE
Comments :
- problem with TAE mode FIXED
- TELL1 firmware updated => THE DETECTOR SPECIFIC HEADER IS NO LONGER IN THE DATA BODY
- The special patch in decoding software is NO LONGER NEEDED !!
- Everything OK but LED data need to be time-aligned
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
hcal_1212_0001 |
xx00 |
??? |
LED |
calibration |
300Hz |
3(TAE) |
100% |
HV:1.3kV - LED OFF - FEB22#14 not activated |
hcal_1212_0002 |
xx00 |
??? |
LED |
calibration |
500Hz |
5(TAE) |
100% |
HV:1.3kV - LED OFF |
hcal_1212_0003 |
xx0x |
23973 |
LED |
calibration |
300Hz |
3(TAE) |
100% |
HV:1.3kV - LED ON - delay 140ns |
hcal_1212_0004 |
xx0x |
81615 |
LED |
calibration |
300Hz |
3(TAE) |
100% |
HV:1.3kV - LED ON - delay 130ns |
hcal_1212_0005 |
xx0x |
26764 |
LED |
calibration |
500Hz |
5(TAE) |
100% |
HV:1.3kV - LED ON - delay 120ns |
hcal_1212_0006 |
xx0x |
72763 |
LED |
calibration |
500Hz |
5(TAE) |
100% |
HV:1.3kV - LED ON - delay 120ns |
THE CALO BANK IN ALL FILES BELOW WERE CREATED WITH TELL1 FIRMWARE VERSION < v2.1
--> THE BANK FORMAT CONTAINS AN EXTRA 4 BYTES HEADER WORD (DETECTOR SPECIFIC WORD)
THE DECODING REQUIRED A PATCHED VERSION OF THE DECODING SOFTWARE
(WITH TOOL.DETECTORSPECIFICWORD = TRUE )*
ALL FILES ABOVE WERE OBTAINED WITH TELL1 F/W VERSION > 2.1
THE STANDARD DECODING TOOL (WITHOUT PATCHING OPTION) CAN BE USED*
Data (07/12/2007)
Description : Hcal C-side commissioning
Data content : LED data - Full Hcal C-side (4 TELL1 - crate 22 & 23) & ODIN-FE
Comments :
- some problems with TAE mode : MEP does not contain consecutive BCIDs but only T0 at consecutive orbit step
- Bunch Counter Reset NOT activated in ODIN setup except for data file hcal_1207_001
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
hcal_1207_0001 |
xx0x |
786 |
LED |
calibration |
112Hz |
3(TAE) |
5% |
HV:1kV Delay:140ns |
hcal_1207_0002 |
xx0x |
1042 |
LED |
calibration |
112Hz |
3(TAE) |
5% |
HV:1kV Delay:140ns |
hcal_1207_0003 |
xx0x |
2162 |
LED |
calibration |
112Hz |
3(TAE) |
50% |
HV:1.3kV Delay:140ns |
hcal_1207_0004 |
xx0x |
1838 |
LED |
calibration |
112Hz |
3(TAE) |
100% |
HV:1.3kV Delay:140+5ns |
hcal_1207_0005 |
xx0x |
1893 |
LED |
calibration |
112Hz |
3(TAE) |
100% |
HV:1.3kV Delay:140+10ns |
hcal_1207_0006 |
xx0x |
1986 |
LED |
calibration |
112Hz |
3(TAE) |
100% |
HV:1.3kV Delay:140+15ns |
hcal_1207_0007 |
xx0x |
167 |
LED |
calibration |
112Hz |
3(TAE) |
100% |
HV:1.3kV Delay:140+20ns |
Data (05/12/2007)
Description : Prs C-side commissioning
Data content : data from Prs crate 0 (FEB 6 removed) -> 1 TELL1 [+ ODIN(07)-FE]
Comments :
- DAQ of FE noise and FE+VFE noise
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
prs_1205_0001 |
x000 |
?? |
noise |
periodic |
11kHz |
1 |
5% |
FE noise |
prs_1205_0001 |
x000 |
?? |
noise |
periodic |
11kHz |
1 |
5% |
FE noise |
prs_1205_0003 |
x000 |
?? |
noise |
periodic |
11kHz |
1 |
5% |
VFE + FE noise |
prs_1205_0004 |
xx00 |
?? |
noise |
calibration |
500 Hz |
3 (TAE) |
5% |
VFE + FE noise |
prs_1205_0005 |
xx00 |
?? |
noise |
calibration |
5 Hz |
3 (TAE) |
100% |
VFE + FE noise |
prs_1205_0006 |
xxxx |
1799 |
noise |
calibration |
5 Hz |
3 (TAE) |
100% |
ODIN-FE bank added |
* Problem : observed with Wireshark that MEPs are periodically lost - seems to be due to a bad configuration of ODIN/TELL1
As a consequence the effective DAQ rate is smaller than quoted in the table
November 2007
Data (26/11/2007)
Description : Hcal C-side commissioning including L0-path (L0calo --> L0DU --> ODIN).
Data content : Full Hcal C-side (4 TELL1 - crate 22 & 23) & ODIN-FE
Comments :
- some problems with TAE mode to be solved (non consecutive BCID in MEP)
- electronics problems (missing FEB data, ADC(ch 31)=0 have been cured
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
hcal_1126_0007 |
xxxx |
222 |
LED |
L0-external |
500Hz |
5 (TAE) |
2% |
|
hcal_1126_0008 |
xxxx |
2522 |
LED |
L0-external |
500Hz |
5 (TAE) |
2% |
hcal_1126_0009 |
x0xx |
11205 |
NO LED |
L0-external |
10Hz |
1 |
100% |
|
hcal_1122_0001 |
xx00 |
?? |
noise |
calibration |
200Hz |
5 (TAE) |
0.5 % |
all night long run with CROC/FEB on --> noise |
hcal_1126_0011 |
xx00 |
?? |
LED |
L0-external |
200Hz |
2 |
2 % |
L0DU bank added |
Data (08/11/2007)
Description : Hcal C-side commissioning (Outer area) including L0-path (L0calo --> L0DU --> ODIN).
Data content : Outer Hcal C-side (2 TELL1 - crate 23) & ODIN-FE
Comments :
- electronics problems : missing FEB data, ADC(ch 31)=0
- problems with TAE mode to be solved (non consecutive BCID in MEP)
Analysis of these data is detailed in the following presentation :
http://indico.cern.ch/getFile.py/access?contribId=3&resId=1&materialId=slides&confId=24065
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
hcal_1108_0007 |
xxxx |
750 |
LED |
L0-external |
500Hz |
5 (TAE) |
1.5% |
|
hcal_1108_0008 |
xxxx |
4101 |
LED |
L0-external |
500Hz |
5 (TAE) |
1.5% |
|
hcal_1108_0008 |
xxxx |
19 |
NO LED |
L0-external |
few/mn |
5 (TAE) |
100 % |
|
September 2007
Data (27/09/2007)
Data content : full HCAL (8 TELL1)
Data : internal TELL1 generator
Comment :
- partially corrupted data format - useful to check the DAQ software behaviour
file index |
mdf/raw/hbook/root |
#sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
hcal_0927_0002 |
x000 |
?? |
generator |
periodic |
xx |
1 |
xx |
partially corrupted data banks |
May 2007
Data (02/05/2007)
Description : First step toward the Hcal commissioning (Inner C-side area) including L0-path (L0calo --> L0DU --> ODIN).
Data content : Inner HCAL C-side (2 TELL1 - crate 22)
Comments :
- electronics problems : missing FEB data, ADC(ch 31)=0
- problems with TAE mode to be solved (non consecutive BCID in MEP)
file index |
mdf/raw/hbook/root |
# sequences |
data |
Triggger type |
rate |
BX sequence |
Storage scaling |
comment |
hcal_0502_0010 |
xx00 |
?? |
LED |
L0-external + calibration |
1.8kHz |
3 |
5 % |
May global commissioning of HCAL |
hcal_0502_0011 |
xx00 |
?? |
LED |
L0-external + calibration |
1.8kHz |
3 |
5 % |
May global commissioning of HCAL |
-- OlivierDeschamps - 29 Nov 2007